# Tree Attention: Topology-Aware Decoding for Long-Context Attention on GPU Clusters

### v3

@@ 4.  Self-Attention as a Gradient
Here, we show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we define a scalar function that depends on the keys, queries, values and additionally on an auxiliary vector that we refer to as the source $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self-attention operation. We need the source in order to write down the generating function of the moments of the distribution above. It is also the variable with respect to which we can Taylor-expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
$\sum^{N}_{a=1}\textrm{softmax}(q\cdot k_{a})v_{a}=\frac{\partial F}{\partial\zeta}\bigg|_{\zeta=0}.$ (5)
@@
This terminology is inspired by work on energy-based models in machine learning Beal (2003); LeCun et al. (2006); Song and Kingma (2021). A summary of variables and indices is provided in appendix  12.
4.1.  The energy function
@@
Here, we show how the energy function is given by the cumunlant-generating function associated to the distribution given by attention scores. Taking inspiration from statistical mechanics, where an analogous cumulant-generating function defines the Helmholtz Free energy Wikipedia (2024), we dub our cumulant-generating function the energy function for self-attention.
@@
$P_{a}=\frac{e^{q\cdot k^{T}_{a}}}{\sum^{N}_{i=1}e^{q\cdot k^{T}_{i}}},$ (7)
and satisfies:
$\sum^{N}_{a=1}P_{a}=1.$ (8)
@@
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We proceed focusing on the single-head case, as it makes the presentation simpler, and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 11.
4.2.  Bayesian interpretation
@@
$\frac{\partial\Gamma}{\partial\zeta_{i,A}}=0,$ (22)
$\frac{\partial\Gamma}{\partial z_{i,A}}=0.$ (23)
@@
5.  Tree Attention
5.1.  Computing the energy efficiently
$F_{dec}=\log\sum^{N}_{a=1}\exp(q\cdot k_{a}^{T}+\zeta\cdot v^{T}_{a})$ (27)
$\equiv\textrm{logsumexp}_{a}(\{q\cdot k^{T}_{a}+\zeta\cdot v^{T}_{a},a=1,\cdots,N\}).$ (28)
$\textrm{logsumexp}_{a}(\{T_{a},\textrm{logsumexp}_{a}(\{R_{a},S_{a}\})\})=$
$\textrm{logsumexp}_{a}(\{\textrm{logsumexp}_{a}(\{T_{a},R_{a}\}),S_{a}\}),$
$\max_{a}(\{\max_{a}(\{T_{a},R_{a}\}),S_{a}\})=$
$=\max_{a}(\{T_{a},\max_{a}(\{R_{a},S_{a}\})\}).$
Theorem 1.
@@ 5.1.  Computing the energy efficiently
Algorithm 1 Single Query Energy Forward
@@
4:Compute $m=\textrm{Reduce}(\max,r)$ by doing a tree reduction.
6:Compute $O=\textrm{Reduce}(\textrm{logsumexp},r)$ by doing a tree reduction.
7:Save $O,m$ for gradient w.r.t $\zeta$.
8:Return $O$
5.2.  Efficient parallel decoding
In our case, we want to compute the gradient of the energy function with respect to $\zeta_{A}$ and then set it to zero. This can be done with automatic differentiation engines having set $\zeta$ to be a tensor of zeros from the very outset. We can however manually implement a gradient with respect to $\zeta$ pass of the above Algorithm  1 that doesn’t materialize $\zeta$ in Algorithm 2 below. Note in particular that when we set $\zeta_{A}=0,\,\,\forall A\in\{1,\cdots,d_{h}\}$ then $O$ involves only the logsumexp of the dot product between queries and keys.
Algorithm 2 Single Query Energy Gradient w.r.t $\zeta$
[…]
2:Scatter a copy of $\mathbf{q},m$ and $O$, and each $\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}$ to each of the $p$ processors.
3:In parallel compute $r_{\hat{a}}=\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{a}}-m$
4:Compute $R_{\hat{a}}=\exp(r_{\hat{a}}-O)\cdot v_{\hat{a}}$
5:Compute $z=\textrm{Reduce}(\textrm{sum},R)$
6:Return $z$
Notice here that by storing $O,m$ for backward, the only remaining reduction operation that needs to be performed is the one in line 5 of the above algorithm. This single reduction takes $O(N/p)$ time to compute the local sums on each device and $\log p$ time to communicate and combine partial results, and therefore we get the same asymptotic complexity as the logsumexp calculation.
Algorithm 3 Tree Decoding
[…]
2:Scatter $\mathbf{q}$ to each GPU.
3:On each GPU, use Flash Attention 2 to compute $\textrm{o}=\frac{\sum_{\hat{a}}\exp(\mathbf{q}\cdot\mathbf{k}_{\hat{a}}^{T})\mathbf{v}_{\hat{a}}}{\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})}$ and $\textrm{lse}=\log\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})$.
4:Recompute the global max $m=\textrm{{Allreduce}}(\textrm{max},\textrm{lse})$.
5:Obtain the local numerator and denominator by then computing: $\textrm{n}=\textrm{o}*\exp(\textrm{lse}-m),\textrm{d}=\exp(\textrm{lse}-m)$.
6:Compute global numerator and denominator $\textrm{n}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{n}),\textrm{d}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{d})$.
7:Return result $z=\frac{\textrm{n}_{g}}{\textrm{d}_{g}}$.
5.3.  Implementation and topology-awareness
Figure 2: NCCL Send/Recv between two H100 GPUs intra-node and inter-node. GPU clusters offer a two-tier topology whereby intra-node bandwidth is significantly higher than inter-node. Algorithms such as Tree Attention exploit this topology by reducing inter-node communication requirements, enabling better overlap of communication with computation.
While the theoretical analysis above indicates that we should see speedups when using tree-based reductions, this is not necessarily guaranteed in practice due to various potential overheads. However, importantly, beyond its asymptotic benefits, Tree Attention benefits from taking advantage of the two-level topology which is standard in modern GPU clusters.
Ring Attention is inherently not topology-aware, and only scales within a network of homogeneous bandwidth. However, this is in conflict with the two-level network topology of modern GPU clusters, which use high-bandwidth interconnects within nodes (NVLINK or PCIe) and comparatively lower-bandwidth interconnects across nodes (InfiniBand or Ethernet). The interconnects greatly differ in bandwidth and latency (see Figure 2). Therefore, Ring Attention is bottlenecked by the slowest interconnect, and cannot always overlap the attention computation with communication. We discuss this point further in 6.3. Tree Attention improves on Ring Attention by using network topology-aware communication patterns to increase overlap of computation and communication, and decrease this scalability bottleneck on communication from the distributed attention computation.
In our empirical experiments , we use Flash Attention 2 Dao (2023) within each device, both for our algorithm and for Ring Attention [^1] A JAX-based Ring Attention implementation that uses Flash Attention 2 can be found here: https://github.com/nshepperd/flash_attn_jax.. We provide a simple JAX implementation of our method in Appendix 9.. Note that our method mirrors Flash Decoding fla (2024) except in that case, the parallelization happens at the level of different streaming multiprocessors (SMs) within a GPU whereas we parallelize between different GPUs. All computations are performed in BF16.
6.  Results
Similar to Ring Attention, Tree Attention is an exact computation of attention. Since training and evaluation metrics are the same as for attention, our experimental results are focused primarily on latency in section 6.1., peak memory usage in section 6.2. and communication volumes in section 6.3.. Since our algorithm computes numerically identical results as the forward pass of standard attention, our performance results transfer seamlessly to transformer architectures.
We performed all experiments on a DGX H100 cluster consisting of 16 nodes, each containing 8 H100 GPUs. All GPUs within the node are connected via an all-to-all NVLINK 4.0 (900GBps) topology. Nodes are connected to each other via 8 InfiniBand NDR interconnects per node (1 per GPU), each of which provides 400 Gbps (leading to an aggregate 3.2 Tbps node injection bandwidth).
6.1.  Latency
Our latency results shows how Tree Attention improves over Ring Attention as we increase the sequence length in Fig. LABEL:fig:relative_times and increase the number of GPUs in Fig. LABEL:fig:absolute_times. As the plots demonstrate, as we scale the sequence length or the number of GPUs, the gap between Tree Attention and Ring Attention execution time widens asymptotically. Remarkably, Tree Attention achieves close $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We expect this trend to continue for larger sequence lengths.
6.2.  Memory cost
(b) Peak memory usage of a single attention block with Tree Attention vs Ring Attention when sharded between two RTX 4090s. Results taken using the JAX memory profiler on one GPU. The difference in peak memory scales with hidden size and sequence length.
[…]
$\textrm{Mem}_{\text{ring}}$ $=4btd+2bd$ (29)
$\textrm{Mem}_{\text{tree}}$ $=2btd+2bd+2bn_{h},$ (30)
6.3.  Communication volume
For Ring Attention’s P2P communication strategy, the total volume of data being communicated between devices (in units of number of tensor elements) per iteration is given by:
$V_{ring}=2btd\times p$ (31)
$V_{\texttt{Allreduce}}=2\times\frac{p-1}{p}\times\textrm{numel}.$ (33)
$\textrm{numel}\left(\textrm{n},\textrm{d},m\right)=bd+2bn_{h}.$ (34)
Note that we first perform on device the local reductions to obtain the local numerator and denominator on each device which consequently makes it so that $t$, i.e. the size of the local sequence chunk doesn’t appear in the above expression. We then obtain:
$V_{Tree}=2\frac{p-1}{p}\times\left(bd+2bn_{h}\right).$ (35)
Our theoretical analysis shows that per iteration our algorithm maintains a lower communication volume than Ring Attention. Note however that Ring Attention when performed in the training setting with many queries overlaps communication and computation so as to hide it’s communication costs. However, overlapping communication and computation in the decoding case is infeasible because of how fast the attention computation on a single GPU is relative to how long it takes to communicate the chunk of keys and values between two devices.
Concretely, let us take the example of decoding from a context of length 640000 split between $8$ GPUs within one node. Let us take a hidden size of 2048 and fix our data type to be bfloat16. Each device for decoding takes $O(10^{-5})$ seconds to perform the Flash Attention computation. The time it takes to move the keys and values of the corresponding size between adjacent GPUs as per Fig. 2 is roughly $O(10^{-3})$ seconds. The latency incurred between nodes is even greater and therefore overlapping isn’t feasible due to this disparity in timescales.
7.  Discussion and Conclusion
In this paper, we have derived the energy function for self-attention and demonstrated how the computation of the derivative of this function provides a novel and efficient method for computing attention in parallel. This advantage is especially apparent when performing decoding across multiple devices, in which case our Tree Attention enables us to substantially outperform Ring Attention with an asymptotically superior algorithm, with $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We also see that the AllReduce operation that we use involves sending partially reduced objects, which greatly reduces the volume of communicated data as well as the peak memory requirement.
Our introduction of a unique energy function for self-attention develops interesting connections between attention and other related models such as Hopfield networks and the general notion of associative memories. Given this energy function, it is possible to mathematically analyze its loss landscape and dynamics, as well as begin to understand how the attention operation could potentially be improved. Additionally, since the energy function can be framed as a variational free energy, it becomes possible to interpret the attention operation through a Bayesian lens and understand the generative model that is implicit in its operation. We hope that these mathematical derivations aid research in understanding the properties of transformer networks, as a deep understanding of the Hopfield energy function has done for Hopfield networks.
Theoretical asymptotic improvements should be demonstrable in the single-device case, but in practice, this advantage is not realized. SMs in H100 and older architectures communicate through shared global memory, not in a peer-to-peer fashion like individual GPUs in a multi-device setup, making the reduced communication negligible. However, recent experimental instructions in the H100 have enabled peer-to-peer SM communication  NVIDIA (2024), suggesting that these instructions could lead to speedups over Flash Attention on a single device.

### v4

@@ C.1 Proof of Observation 1
Here, we show how the self-attention operation can be written as the gradient of an energy function. In particular, we define a scalar function that depends on the keys, queries, values and additionally on an auxiliary vector that we refer to as the source $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self-attention operation. We need the source in order to write down the generating function of the moments of the distribution above. It is also the variable with respect to which we can Taylor-expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
$\sum^{N}_{a=1}\mathrm{softmax}(q\cdot k_{a})v_{a}=\frac{\partial F}{\partial\zeta}\bigg|_{\zeta=0}.$ (15)
@@
This terminology is inspired by work on energy-based models in machine learning Beal (2003); LeCun et al. (2006); Song & Kingma (2021). A summary of variables and indices is provided in appendix  G
@@
We first show how the energy function is given by the cumulant-generating function associated to the distribution given by attention scores. Taking inspiration from statistical mechanics, where an analogous cumulant-generating function defines the Helmholtz Free energy (Landau & Lifshitz, 1958), we dub our cumulant-generating function the energy function for self-attention.
@@
$P_{a}=\frac{e^{q\cdot k^{T}_{a}}}{\sum^{N}_{i=1}e^{q\cdot k^{T}_{i}}}.$ (17)
@@
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We proceed focusing on the single-head case, as it makes the presentation simpler, and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix F
C.2 Bayesian interpretation
@@
$\frac{\partial\Gamma}{\partial\zeta_{i,A}}=0,\frac{\partial\Gamma}{\partial z_{i,A}}=0.$ (31)
@@
C.3 More Performances Results with a LLama Transformer Model
To extend our work in section 6.4, and to demonstrate that Tree Attention can be successfully applied to a range of hardware setups, we also experiment with running Llama3.2-1B on a dual NVIDIA RTX 4090 setup. The two 4090s are connected via PCIe networking. Even in this case, we observe a significant 4x speedup (growing to 5x at longer sequence lengths) of Tree Attention over Ring Attention for autoregressive decoding.
Table 2: Average Decoding Time (in seconds) comparisons with prefill stage using the 1B Llama 3.2 model with Tree Attention (ours) and Ring Attention (SOTA) across various sequence lengths for 4090s. Average results and standard error ($\pm$) are computed using 10 trial runs.
Sequence Length Tree Attention Ring Attention Speedup
Time (s) Time (s)
8000 0.34 $\pm$ 0.05 1.38 $\pm$ 0.07 $\times$4
16000 0.58 $\pm$ 0.07 2.77 $\pm$ 0.04 $\times$5
20000 0.74 $\pm$ 0.01 3.47 $\pm$ 0.04 $\times$5
32000 1.01 $\pm$ 0.02 5.45 $\pm$ 0.03 $\times$5
Appendix D Appendix: JAX code
Below is the tree_flash_decode method. Our full code base is available here: https://anonymous.4open.science/r/tree_attention-7C32.
import jax
from jax import lax
import jax.numpy as jnp
from functools import partial
from jax.sharding import Mesh,NamedSharding, PartitionSpec as P
from jax.experimental import mesh_utils
from jax.experimental.shard_map import shard_map
from flash_attn_jax.flash import _flash_mha_vjp
in_specs=(P(None, None, None, None), P(None, ’i’, None, None), P(None, ’i’, None, None))
out_specs=P(None, None, None)
@jax.jit
@partial(shard_map, mesh=mesh, in_specs=in_specs, out_specs=out_specs, check_rep=False)
def tree_flash_decode(q, k, v):
def flash_num_lse(q, k, v, config=dict(softmax_scale=1.0, is_causal=False, window_size=(-1, -1))):
tup = _flash_mha_vjp.fwd(q, k, v, config)
res,lse = tup[1][3],tup[1][4]
return res,lse
loc_res, loc_lse = flash_num_lse(q, k, v)
a_max_global = lax.pmax(loc_lse, axis_name=’i’)
num_global = lax.psum(loc_res * jnp.exp(loc_lse - a_max_global), axis_name=’i’)
den_global = lax.psum(jnp.exp(loc_lse - a_max_global), axis_name=’i’)
return (num_global / den_global)
The function uses Flash Attention 2 Dao (2023) to compute the local numerator and denominator, both of which are accumulated between devices using an Allreduce (which is what psum and pmax call). NCCL determines in what pattern these results are communicated.
Appendix E Theorem 1 Proof
We prove theorem 1 below.
Proof.
Sequential Case: On a single GPU, the reduction operation over an array of size $N$ has a time complexity of $O(N)$ since the processor must sequentially process each element.
Parallel Processing with $p$ Processors: Divide the array of size $N$ into $p$ chunks, each of size $\frac{N}{p}$. Each processor performs the reduction operation on its chunk independently. The time complexity for each processor is $O\left(\frac{N}{p}\right)$.
Combining Partial Results: The partial results from the $p$ processors need to be combined. Using a tree pattern for reduction, the partial results can be reduced in $O(\log p)$ steps. Each step involves combining pairs of results, halving the number of results at each step until only one result remains.
Total Time Complexity: The total time complexity is the sum of the time complexities for processing the chunks and combining the results:
$O\left(\frac{N}{p}\right)+O(\log p).$
This proves that the time complexity of a reduction involving an associative operation over an array of size $N$ is $O\left(\frac{N}{p}+\log p\right)$ when using $p$ parallel processors, and it reduces to $O(\log N)$ when the number of processors is equal to the size of the array. ∎
Appendix F Computing safe softmax
While, mathematically, attention utilizes the softmax operation, in practice this is often numerically unstable using relatively low precision operations. To address this, a mathematically equivalent function, the ‘safe softmax’ is instead used which subtracts all dot products in the exponential by the max. This ensures that all values being exponentiated are less than 1 and hence less likely to explode and cause numerical instability. Here, we demonstrate that our energy function approach also can account for safe softmax.
Let us suppose we compare our generating function
$F_{tot}=\sum_{i}\log\sum_{a=1}^{i}\exp\left(q_{i}\cdot k_{a}^{T}+\zeta_{a}\cdot v^{T}_{a}\right)$ (35)
and a slightly modified one:
$F^{\prime}_{tot}=\sum_{i}\log\sum_{a=1}^{i}\exp\left(q_{i}\cdot k_{a}^{T}+\zeta_{i}\cdot v^{T}_{a}-m_{i}\right).$ (36)
When we take the derivative of these two quantities, we see that we get the same result:
$\frac{\partial F_{tot}}{\partial\zeta_{i}}\bigg|_{\zeta_{i}=0}=\frac{\partial F^{\prime}_{tot}}{\partial\zeta_{i}}\bigg|_{\zeta_{i}=0}.$ (37)
To see it explicitly:
$\frac{\partial F^{\prime}_{tot}}{\partial\zeta_{i}}\bigg|_{\zeta_{i}=0}=\frac{\sum^{i}_{a=1}\exp(q_{i}\cdot k^{T}_{a}-m_{i})v_{a}}{\sum^{i}_{a=1}\exp(q_{i}\cdot k^{T}_{a}-m_{i})}$ (38)
$\frac{\sum^{i}_{a=1}\exp(q_{i}\cdot k^{T}_{a})v_{a}}{\sum^{i}_{a=1}\exp(q_{i}\cdot k^{T}_{a})}.$
Normally, when computing the softmax in an online fashion, this procedure is performed where $m_{i}$ is the row max of $q\cdot k^{T}$. This shift makes it so that the sum of exponentials doesn’t lead to overflows.
Appendix G Notations for equations
Here is a summary of the various variables and indices that will be used in the coming sections:
TABLE I: Variable names. (39)
@@ Appendix G Notations for equations
TABLE II: Index names and ranges. (40)